OpenStudy (anonymous):

6. Translate each sentence into an equation than find each number. seven more than twice a number is 17

OpenStudy (saifoo.khan):

Suppose the number be: x (7 MORE) than (TWICE the number) is (17) 7 + 2x = 17

OpenStudy (anonymous):

whats the number?

OpenStudy (anonymous):

5

OpenStudy (saifoo.khan):

Solve for x.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

solve for x:\[7+2x=17\]subtract 7 from both sides:\[2x=10\]divide both sides by 2:\[x={10\over2}\] which is.....? \[\large 5~\ddot\smile \]

OpenStudy (anonymous):

twenty more than three times a number is -4

OpenStudy (anonymous):

3x+20=-4 3x=-24 x=-8